OpenStudy (anonymous):

(Medal for correct answer(s)) Use elimination to find the solution to the system of equations. 7x+3y=23 6x-6y= -6 A. x= 0, y= 1 B. x= 2, y= 3 C. x= 3, y= 2/3 D. x= 7, y= -26/3

OpenStudy (anonymous):

Ok I can Help

OpenStudy (anonymous):

There are many ways of elimination. In this elimination you want to solve by adding the functions together or subtraction of 1 from the other to get 1 value left and then using that value to solve 1 equation for the last value left.

OpenStudy (anonymous):

Thanks @Shashank4488 think you can help me with more?

OpenStudy (anonymous):

Use the graph below to estimate the solution to the system of equations shown. A. x= 1/2, y= 3/4 B. x= -1/2, y= 3/4 C. x= -3/4, y=1/2 D. x= 3/4, y= -1/2

OpenStudy (anonymous):

correct! Next question! Use elimination to find the solution to the system of equations. 7x+2y=-63 2x+5y=-18 A. x= -65/7, y= 1 B. x= 4, y= -91/2 C. x= -1, y= -16/5 D. x= -9, y= 0

hero (hero):

@Shashank4488, I deleted your answers, therefore the medals you received have been taken away and you will not receive any credit for just giving the answer. In fact, the only thing that has been been accomplished here (if it can even be called an accomplishment) is that the user has been given the answer without any explanation or understanding of how to arrive at the correct answer on his or her own in similar problems without help. Basically, in this case, no facilitating, guiding, or demonstration of the process of learning in relation to problem solving was given which means, the user didn't learn anything and ultimately, the foundational goals of Open Study have been compromised. Therefore, I recommend that you re-evaluate your approach to helping students here. Otherwise, more appropriate actions will be taken to resolve this issue.

OpenStudy (anonymous):

I Have solved it With Explanation , U didnt see

hero (hero):

You posted mostly answers. Posting the answer before explaining is still considered giving answers. Facilitate, guide, demonstrate. Give the user a chance to express what they believe the answer is after you explain. Responses such as "the answer is B" will never be acceptable here.

OpenStudy (anonymous):

In First question Why did u delete ? I had given with explanation What the hell is this?

hero (hero):

I read everything you did. Once you give an answer for any question, all of your responses will be deleted to take away the medals you did not earn. If you want to earn medals, you'll have to work on your approach to helping students. I will be reviewing your other questions as well.

OpenStudy (anonymous):

I Asked Why Did u delete my first Answer What the hell is this . If I give Answer with no exaplanation then delete that question only Okay I had given exaplanation of first question below

OpenStudy (anonymous):

He wasnt the only one who maybe typed just an answer without explanation

OpenStudy (anonymous):

Your first answer consisted of "1000%" it did not have an explanation. The explanation followed after, I think that was what Hero was referring to

hero (hero):

@PoohDop3, I am aware of that, but I cannot see everything here. Part of the expectation here is that when users see other users giving answers, they are expected to report them using the "report abuse" button.

OpenStudy (anonymous):

Ok I was just saying

hero (hero):

@Shashank4488 you must understand that this site is not intended to be comparable to Yahoo Answers

OpenStudy (anonymous):

U r just seeing me not any other students who just give Answer Only . I just give two or three answer n in rest question I have explained Okay U r just seeing me See other students as well @PoohDop3 U r right

hero (hero):

If you observe other users giving answers, report them. I can't "see" everybody.

OpenStudy (anonymous):

Okay Thanks :) But I had given with exaplanation in first answer It hurted me , U deleted my Answer n snatched my medals , It hurted , Why did u delete that Answer , I want know ?

hero (hero):

@Shashank4488, I had to delete all of your responses to take away the medal. That's the way OS is setup. I'm sorry bro.

OpenStudy (anonymous):

I mean its not that big of a deal to be honest the point is to be sure you include explanation next time thats all its really not that big

OpenStudy (anonymous):

Its Okay brother :)